OpenStudy (anonymous):

since derivative of e^x is e^x then would the derivative of e^2x be e^2x?

OpenStudy (anonymous):

2e^2x

OpenStudy (anonymous):

r u guys talking to me?

OpenStudy (anonymous):

who r u guys talking to?

jimthompson5910 (jim_thompson5910):

you would use the chain rule to get 2*e^(2x)

jimthompson5910 (jim_thompson5910):

basically the derivative of e^(2x) is e^(2x) times the derivative of 2x, which is 2 so e^(2x) * derivative of 2x = e^(2x) * 2 = 2*e^(2x)